OpenStudy (anonymous):

Write AB with A(-7, 5) and B(-3, -7) in component form. 1. <-4, 12> 2. <4, -12> 3. <4, 12> 4. <-4, -12>

OpenStudy (amorfide):

AB= B-A so do the x ordinate of B- x ordinate of A do the same for Y ordinates

OpenStudy (anonymous):

which number is the answer? @amorfide

OpenStudy (amorfide):

Work it out, for example if I had two coordinate points A= (4,2) and B=(2,4) AB=B-A=(2-4,4-2) AB=(-2,2) now do the same for your question

OpenStudy (anonymous):

i got it thanks!

OpenStudy (anonymous):

< 4, -12>

OpenStudy (amorfide):

that is correct